The debt is amortized by the periodic payment shown. Compute​ (a) the number of payments required to amortize the​ debt; (b) the outstanding principal at the time indicated.
Debt Principal $12000
Debt Payment $ 1057
Payment Interval 6 months
Interest Rate 3%
Conversion Period -semi annually
Outstanding Principal​ After: 8th payment
​(a) The number of payments required to amortize the debt is
nothing.
​(Round the final answer up to the nearest whole number. Round all intermediate values to six decimal places as​ needed.)
​(b) The outstanding principal is ​$
nothing.
​(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as​ needed.
Answer #1

Solution :-

Debt Principal = $12,000

Debt Payment = $1,057

Semiannual Interest = 3% / 2 = 1.5%

Now Assume number of Payments = n

$12,000 = $1,057 * PVAF ( 1.5% , n )

$12,000 = $1,057 * [ 1 - ( 1 + 0.015 )-n ] / 0.015

0.1702933 = [ 1 - ( 1 + 0.015 )-n ]

( 1 + 0.015 )n = 1.205245

Take log Both sides

n log( 1.015 ) = log ( 1.205245 )

n = 0.081075394 / 0.006466042

n = 12.54

The number of payments required to amortize the debt is 12.54

(b) The outstanding principal after 8 Payment =

Payment Outstanding = 12.54 - 8 = 4.54

= $1,057 * PVAF ( 1.5% , 4.54 )

= $1,057 * 4.356102277

= $4,604.40

If there is any doubt please ask in comments

Thank you please rate